a cube with edge length of 10 cm is inscribed in a sphere, what is the diameter of the sphere?


1
Expert's answer
2022-05-10T11:48:47-0400

The cube's face is a square. The edge of cube is "10" cm. Thus, the length of the side of the square is "10" cm. The diagonal of the square is equal to "10\sqrt{2}" due to the Pythagorean theorem. The diagonal of the cube is: "\sqrt{10^2+2\cdot10^2}=10\sqrt{3}". It is equal to the diameter of the sphere.

Answer: "10\sqrt{3}" cm.